Compute EPS under all three methods of financing the expansion at $6.1 million in sales (first year) and $10.1 million in sales (last year). (Round your answers to 2 decimal places.)

Delsing Canning Company is considering an expansion of its facilities. Its current income statement is as follows:
 



Sales$5,100,000
Variable costs (50% of sales)
2,550,000
Fixed costs
1,810,000
Earnings before interest and taxes (EBIT)$740,000
Interest (10% cost)
220,000
Earnings before taxes (EBT)$520,000
Tax (35%)
182,000
Earnings after taxes (EAT)$338,000
Shares of common stock
210,000
Earnings per share$1.61

 

The company is currently financed with 50 percent debt and 50 percent equity (common stock, par value of $10). In order to expand the facilities, Mr. Delsing estimates a need for $2.1 million in additional financing. His investment banker has laid out three plans for him to consider:
 

  1. Sell $2.1 million of debt at 9 percent.

  2. Sell $2.1 million of common stock at $15 per share.

  3. Sell $1.05 million of debt at 10 percent and $1.05 million of common stock at $20 per share.

  
 

Variable costs are expected to stay at 50 percent of sales, while fixed expenses will increase to $2,310,000 per year. Delsing is not sure how much this expansion will add to sales, but he estimates that sales will rise by $1 million per year for the next five years.
Delsing is interested in a thorough analysis of his expansion plans and methods of financing.He would like you to analyze the following:

 

The break-even point for operating expenses before and after expansion (in sales dollars). (Enter your answers in dollars not in millions, i.e, $1,234,567.)


 The degree of operating leverage before and after expansion. Assume sales of $5.1 million before expansion and $6.1 million after expansion. Use the formula: DOL = (S − TVC) / (S − TVC − FC). (Round your answers to 2 decimal places.)


The degree of financial leverage before expansion. (Round your answer to 2 decimal places.)
  

The degree of financial leverage for all three methods after expansion. Assume sales of $6.1 million for this question. (Round your answers to 2 decimal places.)


Compute EPS under all three methods of financing the expansion at $6.1 million in sales (first year) and $10.1 million in sales (last year). (Round your answers to 2 decimal places.)
Answer #1

a.

Break-even Point
Before Expansion $ 3,620,000
After Expansion $ 4,620,000

Break-even Point = Fixed Costs / Contribution Margin Ratio

b.

Operating Leverage
Before Expansion 3.45 x
After Expansion 4.12 x

Before Expansion : $ 5,100,000 x 50 % / $ ( 2,550,000 - 1,810,000 ) = $ 2,550,000 / $ 740,000 = 3.45

After Expansion : $ 6,100,000 x 50 % / $ ( 3,050,000 - 2,310,000 ) = $ 3,050,000 / $ 740,000 = 4.12

c-1. Degree of financial leverage : EBIT / ( EBIT - Interest )

Degree of Financial Leverage 1.42 x

c-2 .

Degree of Financial Leverage
100 % Debt 2.24 x
100 % Equity 1.42 x
50 % DEbt and 50% Equity 1.78 x

Computatio of DFL after expansion under three methods of financing :

Method 1. : DFL = $ ( 6,100,000 x 50 % ) - $ 2,310,000 / $ ( 740,000 - 409,000) = 2.24 x

Method 2 : DFL = $ ( 6,100,000 x 50 % ) - $ 2,310,000 / $ ( 740,000 - 220,000) = 1.42 x

meethod 3 : DFL = $ ( 6,100,000 x 50 %) - $ 2,310,000 / $ ( 740,000 - 325,000) = 1.78 x

c-3.

EPS, First Year EPS, Last Year
100 % Debt $ 1.02 $ 7.22
100 % Equity $ 0.97 $ 4.68
50 $ Debt and 50 % Equity $ 1.03 $ 5.98

Computation of EPS :

100 % Debt 100 % Equity 50 % Debt and 50 % Equity
Sales, First Year $ 6,100,000 $ 6,100,000 $ 6,100,000
Variable Cost 3,050,000 3,050,000 3,050,000
Contribution Margin 3,050,000 3,050,000 3,050,000
Fixed Costs 2,310,000 2,310,000 2,310,000
EBIT 740,000 740,000 740,000
Interest 409,000 220,000 325,000
Earnings before taxes 331,000 520,000 415,000
Earnings after taxes 215,150 338,000 269,750
Common shares outstanding 210,000 350,000 262,500
Earnings per Share $ 1.02 $ 0.97 $ 1.03